# End-to-End Autonomous Driving through V2X Cooperation

### v2

@@ Appendix 0.A Experiment Results on More Dataset
No Fusion 0.98 1.79 2.65 1.81 2.0 2.25 1.75 2.0
UniV2X 0.97 1.77 2.66 1.80 1.75 2.0 1.5 1.75
[…]
Table 9: Agent Perception Evaluation Results on V2X-Sim Dataset [25].
method mAP (%$\uparrow$ AMOTA (%$\uparrow$
No Fusion 7.0 0.6
UniV2X 7.6 1.7
[…]
Table 10: Occupancy Evaluation Results on V2X-Sim Dataset [25].
method IoU-n (%$\uparrow$ IoU-f (%$\uparrow$
No Fusion 15.9 18.9
UniV2X 16.1 19.7
Appendix 0.B Ego Status in Planning Performance.
The significance of ego status, particularly its velocity, has been emphasized in discussions on open-loop end-to-end autonomous driving. Velocity aids in planning trajectories akin to those of expert drivers, thereby substantially mitigating L2 Error [66, 29]. In this section, we delve into the impact of ego-vehicle velocity.
0.B.0.1 Experiment Settings.
We conduct two additional experiments:
  • CooperNaut [9] with ego-vehicle velocity: In this experiment, we incorporate ego-vehicle velocity into an 8-dimensional feature and concatenate it with the fused Bird’s Eye View (BEV) feature. The resulting concatenated feature is then passed through a Multi-Layer Perceptron (MLP) to generate the final planning path. We subsequently train this model using the same configuration as the CooperNaut baseline, denoting it as "CooperNaut+V".
  • UniV2X with ego-vehicle velocity: In this experiment, we integrate velocity information into an 8-dimensional feature vector, which is then concatenated with the ego planning query. We then train this model under the same conditions as UniV2X, naming it "UniV2X+V".
0.B.0.2 Experiment Results.
As illustrated in Table 11, our UniV2X model, when augmented with embedded ego-vehicle velocity, demonstrates reductions of 1.94 meters and 0.24% in average L2 error and average collision rate, respectively. Similarly, CooperNaut with ego-vehicle velocity exhibits decreases of 3.36 meters and 0.98% in average L2 error and average collision rate, respectively. In summary, UniV2X achieves superior collision rate results compared to both CooperNaut and CooperNaut+V, reaching state-of-the-art (SOTA) performance in collision rate, regardless of whether ego-vehicle velocity is considered. However, it’s crucial to acknowledge that integrating ego-vehicle velocity into the model may introduce bias, as velocity encapsulates historical information about the ego-vehicle. This inclusion may inadvertently transform the planning problem into one of ego-vehicle trajectory prediction, potentially skewing overall evaluation results. This consideration aligns with research discussed in [66, 29].
Table 11: More Planning Evaluation Results Considering Ego-Vehicle Velocity.
[…]
CooperNaut [9] 3.83 5.26 6.69 5.26 0.59 1.92 1.63 1.38 8.19×$\times 10^{7}$
CooperNaut+V [9] 1.55 1.90 2.26 1.90 0.15 0.15 0.89 0.40 8.19×$\times 10^{7}$
UniV2X 2.60 3.34 4.36 3.43 0.00 0.74 0.74 0.49 8.09×$\times 10^{5}$
UniV2X+V 1.35 1.41 1.71 1.49 0.00 0.15 0.59 0.25 8.09×$\times 10^{5}$
Appendix 0.C Flow Prediction for Latency Compensation
0.C.0.1 Experiment Settings.
[…]
0.C.0.2 Experiment Results.
@@ Appendix 0.C Flow Prediction for Latency Compensation
UniV2X 500 28.5 (-1.0) 23.5 (-0.4)
Appendix 0.D Different Communication Bandwidth Constraints
@@
Appendix 0.E Evaluation Metrics
0.E.0.1 Object Detection.
We evaluate the detection performance using the mean Average Precision (mAP) metric, which is commonly used in previous works such as [60, 3]. The AP is calculated based on the 11-points interpolated precision-recall curve and is defined as follows:
0.E.0.2 Multi-object Tracking.
AMOTA and AMOTP, commonly 3D tracking evaluation metrics [3], are used to measure the performance of multi-object tracking.
0.E.0.3 Online Mapping.
0.E.0.4 Occupancy Prediction.
In line with UniAD [16], we assess the accuracy of predicted occupancy using a methodology similar to that outlined in [15, 67]. Specifically, the Intersection over Union (IoU) metric quantifies categorical segmentations of the entire scene in an instance-agnostic manner. This metric is computed within Bird’s Eye View (BEV) ranges, with evaluations conducted for both near (30m×$\times$30m) and far (50m×$\times$50m) distances.
0.E.0.5 Planning.
We utilize the L2 error and collision rate metrics at different timestamps to assess the planning performance. We recognize the significance of considering the lane violation rate as an additional metric for measuring planning performance, and we plan to incorporate this metric in our future evaluations.
0.E.0.6 Transmission Cost.
Appendix 0.F Qualitative Visualization
@@
(a) Visualization Example: Turn Left.
(b) Visualization Example: Keep Forward.
(c) Visualization Example: Turn Right.
Figure 5: Planning Visualization of UniV2X on DAIR-V2X [60]. UniV2X consistently demonstrates the ability to generate high-quality planning outputs for diverse driving scenarios, including left turns, straight keeping, and right turns.

### v3

@@ Appendix E Ablation Study for Flow Prediction
Experiment Results.
@@
UniV2X 500 28.5 23.5
Appendix F Ablation Study for Communication Constraints
@@
Appendix G Experiments on More Dataset
In this section, we apply UniV2X to the V2X-Sim dataset (Li et al. 2022a) to showcase its effectiveness in V2V scenario. Due to the challenges in converting the dataset map to our format, we have not implemented online mapping and lane fusion in our analysis. Consequently, we are unable to report results for the off-road rate metric.
V2X-Sim Dataset.
The V2X-Sim dataset (Li et al. 2022a) is a synthetic dataset for cooperative autonomous driving, developed through the integration of Carla Simulator (Dosovitskiy et al. 2017) and SUMO (Lopez et al. 2018) co-simulation. This dataset comprises 100 scenes, and each scene contains synchronized image and point cloud sequences, at a key frame rate of 5Hz, from roadside sensors and sensors mounted on five autonomous driving vehicles.
Experiment Settings and Results.
We choose two autonomous vehicles from V2X-Sim and designate one as the ego vehicle to establish V2V scenes. To maintain consistency with the implementation on DAIR-V2X (Yu et al. 2022), we equip each autonomous vehicle solely with a front camera sensor. For this experiment, we utilize only one-fifth of the dataset to streamline the results. Considering the absence of evaluation metrics reflecting online mapping’s impact on planning, such as lane violation, we exclude the online mapping task from this experiment. We proceed with training and evaluation of UniV2X on V2X-Sim using settings akin to those in DAIR-V2X (Yu et al. 2022). Experimental results are detailed in Table 11,  12, and  13.
Table 11: Planning Evaluation Results on V2X-Sim (Li et al. 2022a).
Method L2 Error ($m$$\downarrow$ Col. Rate (%$\downarrow$
[…]
No Fusion 0.98 1.79 2.65 1.81 2.0 2.25 1.75 2.0
UniV2X 0.97 1.77 2.66 1.80 1.75 2.0 1.5 1.75
Table 12: Agent Perception Evaluation Results on V2X-Sim (Li et al. 2022a).
method mAP (%$\uparrow$ AMOTA (%$\uparrow$
No Fusion 7.0 0.6
UniV2X 7.6 1.7
Table 13: Occupancy Evaluation Results on V2X-Sim (Li et al. 2022a).
method IoU-n (%$\uparrow$ IoU-f (%$\uparrow$
No Fusion 15.9 18.9
UniV2X 16.1 19.7
Appendix H Qualitative Visualization
@@
Figure 5: Visualization Example: Turn Left.
Figure 6: Visualization Example: Keep Forward.
Figure 7: Visualization Example: Turn Right.
Figure 8: Planning Visualization of UniV2X on DAIR-V2X (Yu et al. 2022). UniV2X consistently demonstrates the ability to generate high-quality planning outputs for diverse driving scenarios, including left turns, straight keeping, and right turns.